the cook time of a chicken depends on the weight of the chicken as given by the following formula: C=15+50w where C is the cook time (in minutes) and w is the weight in kilograms.
A) Transpose the formula, Solving for w

B) if a chicken takes 1hr and 35mins to cook, what is the weight of the chicken?

Answer by Theo(13342)

formula is c = 15 + 50w
c is the cook time in minutes
w is the weight in kilograms
subtract 15 from both sides of the equation to get:
c - 15 = 50w
divide both sides of the equation by 50 to get:
(c-15)/50 = w
switch sides of the equation to get:
w = (c-15)/50
the cook time is 1 hour and 35 minutes = 95 minutes.
formula becomes w = (95-15)/50
solve for w to get w = 1.6 kilograms
confirm by going back to the original equation and replacing w with 1.5
you get c = 15 + 50w becomes c = 15 + 50 * 1.6 which becomes c = 15 + 80 which becomes c = 95 minutes.
solution of w = 1.6 kilograms is good.